OpenStudy (anonymous):

The speed of a passenger train is 16mph faster than the speed of a freight train. The passenger train travels 300 miles in the same time it takes the freight train to travel 220 miles. Find the speed of each train.

OpenStudy (saifoo.khan):

\[Speed = \frac{distance}{time}\] \[time = \frac{distance}{speed}\]

OpenStudy (saifoo.khan):

Freight train, 220/x passenger train, 300/x+16

OpenStudy (saifoo.khan):

\[\frac{200}{x}=\frac{300}{x+16}\]

OpenStudy (anonymous):

I was trying to work on this problem and your third response is as far as I got. I don't know how to do the rest. Thank you.

OpenStudy (saifoo.khan):

\[300x = 200x + 3200\] \[x = 32\]

OpenStudy (saifoo.khan):

where are u stuck now? i can explain..

OpenStudy (anonymous):

I came up with the equations, 300=(r+16)t, 220=(r)t. I just used r as the variable. I don't know what to do next to get the speeds for each train.

OpenStudy (saifoo.khan):

Now as the question is saying that they took the same time.. so equal them with respect to time.

OpenStudy (saifoo.khan):

Like in the second thing i wrote was the equation with respect to time.. As both took the same time, so we can equal them to find the variable.

OpenStudy (saifoo.khan):

getting it?

OpenStudy (anonymous):

Yes, I see.

OpenStudy (anonymous):

So are they both 32mph?

OpenStudy (anonymous):

@ saifoo.khan The solution to the following is x = 44, not 32.\[\frac{300}{x+16}=\frac{220}{x}\]

OpenStudy (saifoo.khan):

Oh, Right. LOL.

OpenStudy (saifoo.khan):

freight = x = 44. passenger = x+16 = 60

OpenStudy (anonymous):

Ah, I'm confused now hahah ;p jk

OpenStudy (saifoo.khan):

Freight train, 220/x passenger train, 300/x+16 \[\frac{220}{x}=\frac{300}{x+16}\] \[220x + 3520=300x\] \[x = 44\]

OpenStudy (anonymous):

Thank you for the medals.

OpenStudy (saifoo.khan):

so, freight = x = 44 passanger = x+16 = 44+16 = 60

OpenStudy (saifoo.khan):

Sorry for the typos. THanks Robo! :D

OpenStudy (saifoo.khan):

Still confused? :(

OpenStudy (anonymous):

No, I got it. Thank you! I am just working on more problems like this and left the window open. :)

OpenStudy (saifoo.khan):

Ohh! okay!! cool. if u r stuck, u can ask problems here!!

OpenStudy (anonymous):

Will do! :)

OpenStudy (saifoo.khan):

Cool. [:

OpenStudy (anonymous):

A particular hybrid car travels 162 mi on 3 gal of gas. Find the amount of gas required to make an 810-mi trip. (I am having trouble crossing the products to solve the proportion.)

OpenStudy (saifoo.khan):

that's easy.

OpenStudy (saifoo.khan):

15 gal.

OpenStudy (anonymous):

Wow! I admire your brain. :)

OpenStudy (saifoo.khan):

Thanks!! ^_^ im 16. :D

OpenStudy (anonymous):

Awesome. I hope you get very far. I've been out of school for too long and am just remembering these things. I was your age 10 years ago haha. You're a great help.

OpenStudy (saifoo.khan):

lol, thank you!! Thankyou. this is mathlab right? i help many mathlabbers! :D

OpenStudy (anonymous):

Yes, it is. I've been getting awesome grades for every class... but math... eeh not so great. This website is amazing. I've only got another 9 questions to finish and then i'm done with this checkpoint.

OpenStudy (saifoo.khan):

if u want help, i can help u through! ^_^.

OpenStudy (anonymous):

Okay! Hat sizes are determined by measuring the circumference of one's head in either inches or centimeters. Use ratio and proportion to complete the following table: Hat size: Head circumference: (to nearest 1/5 inch) Head circ. to nearest cm 7 3/8 23 1/5 59 7 1/4 ___ ___ (For the first answer box, round to the nearest 5th inch as needed. For the 2nd box round to the nearest cm as needed.)

OpenStudy (saifoo.khan):

wow, that was long. LOL.

OpenStudy (anonymous):

hahaha. I know! -_-

OpenStudy (saifoo.khan):

Okay! Hat sizes are determined by measuring the circumference of one's head in either inches or centimeters. Use ratio and proportion to complete the following table: Hat size: Head circumference: (to nearest 1/5 inch) Head circ. to nearest cm 7 3/8 23 1/5 59 7 1/4 22.8 ___ (For the first answer box, round to the nearest 5th inch as needed. For the 2nd box round to the nearest cm as needed.)

OpenStudy (saifoo.khan):

Okay! Hat sizes are determined by measuring the circumference of one's head in either inches or centimeters. Use ratio and proportion to complete the following table: Hat size: Head circumference: (to nearest 1/5 inch) Head circ. to nearest cm 7 3/8 23 1/5 59 7 1/4 22.8 58 (For the first answer box, round to the nearest 5th inch as needed. For the 2nd box round to the nearest cm as needed.)

OpenStudy (anonymous):

Woot! Next one... A sample of 141 firecrackers contained 7 duds. How many duds would you expect in a sample of 1692 firecrackers?

OpenStudy (saifoo.khan):

84

OpenStudy (anonymous):

I need to attach an image for the next one since there are two triangles.

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

Ok. :)

OpenStudy (saifoo.khan):

not 9.

OpenStudy (saifoo.khan):

17.25

OpenStudy (anonymous):

It said 17.25 was not correct.

OpenStudy (saifoo.khan):

92/3 ?

OpenStudy (anonymous):

Correct. Next one... Find the variation constant and an equation of variation where y varies directly as x and y = 24 when x=4 The variation constant is k = ?

OpenStudy (saifoo.khan):

6

OpenStudy (saifoo.khan):

b) 6x

OpenStudy (anonymous):

Awesome. Posting another attachment.

OpenStudy (saifoo.khan):

2952?

OpenStudy (anonymous):

It was incorrect. I attached the hint it mentions.

OpenStudy (saifoo.khan):

i hate this one! always get wrong! >_<

OpenStudy (anonymous):

I think it might be the result of 2952*1,166,000???

OpenStudy (saifoo.khan):

NOWAY.

OpenStudy (saifoo.khan):

didn't it chaged the values?

OpenStudy (saifoo.khan):

after u got it wrong?

OpenStudy (saifoo.khan):

i have just 8 more minutes! :O

OpenStudy (anonymous):

Here's an example... Oh noes!

OpenStudy (saifoo.khan):

i will be back after 1 hour them!

OpenStudy (saifoo.khan):

460,570,000

OpenStudy (anonymous):

Okay. I will still have time to submit since I am in a pacific time zone. Correct! This leaves me at 4 more. :)

OpenStudy (saifoo.khan):

let, roll.

OpenStudy (anonymous):

Dun dun dun!

OpenStudy (saifoo.khan):

Hey im back! :D

OpenStudy (saifoo.khan):

just a sec.

OpenStudy (anonymous):

Welcome back, buddy! Sure. :)

OpenStudy (saifoo.khan):

a) k = 140. b) 140/x

OpenStudy (saifoo.khan):

Q21) 64

OpenStudy (anonymous):

Cool! Did you see number 22? I knew it was an inverse variation but I didn't get the 2nd part of the question.

OpenStudy (saifoo.khan):

i wrote it's answer as well, it's 140/x

OpenStudy (anonymous):

D'oh! I meant 23. Silly me. ;p

OpenStudy (saifoo.khan):

1.5

OpenStudy (anonymous):

You probably do these all in your head??? :D

OpenStudy (saifoo.khan):

haha, yea.. i have done the whole Algebra course like 4 times. LOL.

OpenStudy (saifoo.khan):

And final weeks like 7times.

OpenStudy (anonymous):

Oh, snap. I think the only thing that changes in Math Lab are the numbers. I've found a lot of examples online. Some I just don't understand, but I'm getting there. Here's the last one. Weeeeeeeeeee!

OpenStudy (saifoo.khan):

Yea, they know that people can cheat with those answers. Like in Appendixes.

OpenStudy (anonymous):

You will get very far in life... It seems like you help people for the sake of helping them because it's what you enjoy doing. Karma is always good to people like you. :) I have actually been trying to learn this time around because knowledge will help you know what you're talking about. I guess math can be hard if you don't want to learn it, which most people don't.

OpenStudy (saifoo.khan):

3200/9

OpenStudy (saifoo.khan):

hehe, thank you.. thanks alot for the appreciation.. Yes, i enjoy alot and i learn alot. and people get advantage from it.. which is really good. well, there are other reasons as well.. Yes, you are right. I have a pretty good command on math.. so i try helping people out.. since the material provided by UoP people isn't that good..

OpenStudy (anonymous):

I kept getting unresponsive script messages for a few minutes there. And you're welcome. My hands are tied with so many responsibilities that it makes schoolwork nearly impossible. Yeah, some of the material seems pretty outdated so I go on Youtube to try and find the most informational tutorials. They really help. I'm probably going to be back on here tomorrow if I need help with the quiz that is due Sunday.

OpenStudy (saifoo.khan):

Ohh! yes, youtube provides alot of help. in which week are u in? how many questions does it have?

OpenStudy (anonymous):

Just finishing up Week 4. The quiz has 23 questions of weeks 3 and 4. I can do the problems as far as simplifying, but I am not very good at worded problems.

OpenStudy (saifoo.khan):

Alright, but how u gonna tell me that u r online? do u have email fb or something?

OpenStudy (anonymous):

I see your email is posted on your openstudy profile. I will email you from my gmail account.